# Fast Crystal Tensor Property Prediction: A General O(3)-Equivariant Framework Based on Polar Decomposition

### v2

@@ 3.1 The Rationale of Our Framework
where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ (Geiger & Smidt, 2022), and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics. When transmitting rotational information for higher-order tensors like the elastic tensor, the rotation order $\lambda$ must be set to 4 (Yan et al., 2024b). In comparison, for transmitting lower-order tensor rotational information (where $\lambda$ is set to 1 (Yan et al., 2024a)), the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e_{ji}}})$ increases fourfold, thus, the corresponding computational cost increases. If we can transmit the rotational information for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, this would significantly reduce the computational costs.
To this end, we propose the GoeCTP method, which achieves $O(3)$ equivariance without requiring specialized network architecture design. For instance, it enables the transmission of rotational information for higher-order tensors using $\lambda=1$ rather than $\lambda=4$, thereby simplifying the model complexity while maintaining equivariance. GoeCTP leverages polar decomposition, a mathematical technique with significant geometric properties. Polar decomposition decomposes a matrix into two components: an orthogonal matrix, representing a rotation or reflection, and a positive semi-definite matrix (Higham, 1986). The orthogonal matrix captures an $O(3)$-group transformation, while the positive semi-definite matrix represents a stretching or scaling. When applied to crystal structures, polar decomposition effectively separates the spatial directional information of the crystal via the orthogonal matrix, allowing us to directly transfer equivariance without the need for intricate model design. Furthermore, polar decomposition is only performed during the data preprocessing stage and does not occur during network training, thereby introducing virtually no additional computational burden.
@@ 3.2 Our Proposed Framework: GoeCTP
In what follows, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, focusing on how to obtain a standardized crystal position for a crystal with arbitrary spatial direction using polar decomposition. Then, we will describe how the input crystal data is processed and introduce the property prediction network of GoeCTP. Finally, we will explain how proposed GoeCTP achieves equivariant property predictions. An overview of the GoeCTP framework is illustrated in Fig. 2.
R&R Module. The primary function of the R&R Module is to rotate and reflect the crystal to a standardized crystal position in space. In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for the R&R module to perform rotation and reflection operations on the crystal data. Therefore, R&R module adopts the fractional coordinate system to represent the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. This choice simplifies the implementation of the R&R module. Specifically, the application of $O(3)$-group transformations is achieved by applying polar decomposition as follows.
Proposition 3.1
@@
(Polar Decomposition (Hall & Hall, 2013; Higham, 1986; Jiao et al., 2024).) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\exp(\mathbf{S})$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{S}\in\mathbb{R}^{3\times 3}$ is a symmetric matrix and $\exp(\mathbf{S})=\sum_{n=0}^{\infty}\frac{\mathbf{S}^{n}}{n!}$ defines the exponential mapping of $\mathbf{S}$.
Building on the above proposition, it is evident that for a crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, the invertible lattice matrix $\mathbf{L}$ can be uniquely represented by a symmetric matrix $\mathbf{S}$. By defining $\mathbf{H}=\exp(\mathbf{S})$, the lattice matrix $\mathbf{L}$ can also be uniquely represented by a symmetric matrix denoted as $\mathbf{H}$, i.e. $\mathbf{L}=\mathbf{Q}\mathbf{H}$. Furthermore, $\mathbf{H}$ remains invariant with the $O(3)$ group transformation.
Proof 3.1.1
$\mathbf{L}=\mathbf{Q}\mathbf{H}$ (4)
$\mathbf{Q^{\prime}}\mathbf{L}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (5)
By combining Eq. 4 and 5, we can have the following equation:
$\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}.$ (6)
Because $\mathbf{Q^{\prime}}\mathbf{Q}(\mathbf{Q^{\prime}}\mathbf{Q})^{T}=\mathbf{Q^{\prime}}\mathbf{Q}\mathbf{Q}^{T}\mathbf{Q^{\prime}}^{T}=\mathbf{I}$, it follows that $\mathbf{Q^{\prime}}\mathbf{Q}$ is an orthogonal matrix. According to the uniqueness of polar decomposition, we can have $\mathbf{Q^{\prime}}\mathbf{Q}=\mathbf{Q_{1}}$, which implies $\mathbf{Q_{1}}\mathbf{H}=\mathbf{Q_{1}}\mathbf{H_{1}}$. If $\mathbf{H_{1}}\neq\mathbf{H}$, this would violate the uniqueness property of polar decomposition. Therefore, we conclude that $\mathbf{H_{1}}=\mathbf{H}$, indicating that $\mathbf{H}$ remains unaffected by the $O(3)$ group transformation.
In essence, regardless of the $O(3)$ group transformation applied to the crystal in space, the corresponding $\mathbf{H}$ associated with $\mathbf{L}$ remains invariant. Consequently, $\mathbf{H}$ serves as a fixed standardized crystal position for the crystal $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$, allowing the crystal with different position to be consistently rotated to a same position through polar decomposition. The concept of the standardized crystal position can be formally stated as the following proposition :
Proposition 3.2
(Standardized Crystal Position.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$, the standardized crystal position $\mathbf{H}$ is obtained. When an arbitrary $O(3)$ group transformation $\mathbf{Q^{\prime}}$ is applied to a crystal structure, $\mathbf{H}$ remains unchanged.
Therefore, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ formulated as:
$f_{p}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),$ (7)
where $f_{p}(\mathbf{M})$ represents the application of polar decomposition to the lattice matrix $\mathbf{L}$ of $\mathbf{M}$.
@@
As shown in Fig. 2, the crystal $(\mathbf{A},\mathbf{F},\mathbf{H})$ is passed to the subsequent crystal graph construction module for further processing. The orthogonal matrix $\mathbf{Q}$ obtained during this decomposition is passed to the reverse R&R module, ensuring the equivariant transformation of the output tensor properties. Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into a standardized spatial position that is invariant under $O(3)$ space group transformations. With this particular module, the equivariance can be captured, meaning that the subsequent components of GeoCTP are no longer required to account for equivariance.
Crystal Graph Construction. To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data. Here, we use the crystal graph construction from Yan et al. (2024a; b) to describe the structure and relationships within crystals. Specifically, we first convert the fractional coordinate system $(\mathbf{A},\mathbf{F},\mathbf{H})$ into the Cartesian coordinate system $(\mathbf{A},\mathbf{P},\mathbf{H})$ as introduced in Sec. 2. We assume that the output crystal graph is represented as $\mathcal{G}(\mathcal{V},\mathcal{E})$, $\mathcal{V}$ denotes the set of nodes $v_{i}$ in the crystal graph, where each node $v_{i}$ contains atomic features $\mathbf{v}_{i}=(\bm{a}_{i},\hat{\bm{p}}_{i})$. $\mathcal{E}$ represents the set of edges denoted as $e_{ij}$, which are typically constructed based on the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$. When the Euclidean distance $d_{ij}$ between nodes $v_{i}$ and $v_{j}$ is less than a given radius $R$, i.e. $d_{ij}=||\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}||_{2}\leq R$, an edge $e_{ij}$ will be built with edge feature $\mathbf{e}_{ij}=\bm{p}_{j}+k_{1}\bm{l}_{1}+k_{2}\bm{l}_{2}+k_{3}\bm{l}_{3}-\bm{p}_{i}$. Here, $R$ is established based on the distance to the $k$-th nearest neighbor, and different values of $\bm{k}=[k_{1},k_{2},k_{3}]\in\mathbb{Z}^{3}$ represent different edges between nodes $v_{i}$ and $v_{j}$. Since there are no equivariance requirements for subsequent models, any other graph construction methods can be used to replace this part, such as Wang et al. (2024a); Yan et al. (2024a).
@@
Node and Edge Feature Embedding. Building on previous work (Xie & Grossman, 2018; Yan et al., 2024b; a), node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$. Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$. The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding (Lin et al., 2023). Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
@@
Prediction Module. Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al. (2024a); Taniai et al. (2024); Yan et al. (2022); Lee et al. (2024); Wang et al. (2024a), among others. For better performance, we select eComformer (Yan et al., 2024a), which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module. A detailed explanation of eComFormer can be found in Appendix A.1. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows:
$\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}.$ (8)
@@
$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}}),$ (9)
@@
$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{T}.$ (11)
For predicting the higher-order elastic tensor, the conversion process becomes more complex. Specifically, the conversion is expressed as follows:
$C_{ijkl}^{\mathrm{final}}=\sum_{m=1}^{3}\sum_{n=1}^{3}\sum_{p=1}^{3}\sum_{q=1}^{3}\mathbf{Q}_{im}\mathbf{Q}_{jn}\mathbf{Q}_{kp}\mathbf{Q}_{lq}C_{mnpq},$ (12)
where the subscripts denote the specific positions of the elements within the tensor data.
4 Related work
GNN-Based Methods. CGCNN is a pioneering GNN model specifically designed for handling crystal structures (Xie & Grossman, 2018). This model proposed to represent crystal structures as multi-edge crystal graphs. It was applied to predict various single-value properties such as formation energy and band gap. Since then, several GNN methods have been developed to improve upon CGCNN through exploring various network designs or leveraging prior knowledge (Chen et al., 2019; Louis et al., 2020; Choudhary & DeCost, 2021; Das et al., 2023; Lin et al., 2023). These GNN methods are primarily designed for single-value property prediction and do not address the prediction of high-order tensor properties, such as dielectric or elastic tensors. Furthermore, they lack the ability to preserve the equivariance required for accurate high-order tensor property prediction. In contrast, recent studies attempted to ensure equivariance through specialized network architectures (Mao et al., 2024; Lou & Ganose, 2024; Heilman et al., 2024; Wen et al., 2024; Pakornchote et al., 2023; Yan et al., 2024b; Zhong et al., 2023). These approaches generally employ harmonic decomposition to achieve equivariance for tensor properties. Within these network architectures, many operations are required, such as tensor products and combining irreducible representations. These processes significantly increase computational costs, especially when handling higher-order data.
Transformer-Based Methods. Transformers, with their self-attention mechanism and parallel processing capabilities, are particularly well-suited for predicting crystal material properties. Matformer (Yan et al., 2022), one of the earliest Transformer-based networks used for crystal material property prediction, encoded crystal periodic patterns by using the geometric distances between the same atoms in neighboring unit cells. This addressed the issue in earlier GNN-based methods, including CGCNN, MEGNet, GATGNN, and others, which neglected the periodic patterns of infinite crystal structures. Subsequently, more advanced Transformer-based approaches were proposed. Some methods, such as ComFormer (Yan et al., 2024a), CrystalFormer (Wang et al., 2024a), and CrystalFormer (Wang et al., 2024a), typically incorporate either enhanced graph construction techniques or physical priors, exhibiting impressive results in the prediction of single-value properties. Furthermore, DOSTransformer (Lee et al., 2024) is tailored for the density of states prediction, utilizing prompt-guided multi-modal transformer architecture to achieve super performance. Nevertheless, these models are not directly applicable for accurate high-order tensor property prediction due to their inability to capture the necessary equivariance.

### v3

@@ 3.1 The Motivation of Our Framework
where $\bm{f}_{i}^{l}$ is node $i$’s features, $|\mathcal{N}_{i}|$ is the number of neighbors of node $i$, $\mathbf{TP}_{\lambda}$ refers to tensor product layers with rotation order $\lambda$ [7], and $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ represents the equivariant edge feature embedding, which is embedded by the corresponding spherical harmonics. To achieve $O(3)$ equivariance of the model and ensure high prediction accuracy for high-order tensor prediction tasks, such as elastic tensor prediction, the rotation order $\lambda$ is set to 3 [51], i.e., $\mathbf{Y}_{3}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{7}$. In contrast, for lower-order tensor prediction tasks, setting $\lambda$ to 1 is sufficient to achieve the model’s $O(3)$ equivariance, i.e., $\mathbf{Y}_{1}(\hat{\mathbf{e}}_{ji})\in\mathbb{R}^{3}$. Compared to lower-order tensor prediction, the dimension of $\mathbf{Y}_{\lambda}(\hat{\mathbf{e}}_{ji})$ nearly doubles for high-order tensor predictions, leading to a significant increase in computational complexity. Consequently, this model architecture introduces tremendous computational costs for tensor property predictions. If we can achieve the model’s $O(3)$ equivariance and maintain the model performance for higher-order tensors with a lower value of $\lambda$, e.g., $\lambda=1$, or even without relying on harmonic decomposition, computational costs could be drastically reduced. Inspired by the recent exploration of using canonicalization in efficient learning with symmetry, we aim to propose a principled approach based on canonicalization to enable $O(3)$-equivariant tensor predictions for crystal materials.
3.2 $O(3)$ Equivariant Tensor Prediction from the Perspective of Canonicalization
To mathematically describe $O(3)$-equivariant tensor prediction from the perspective of canonicalization, we first introduce and extend several fundamental concepts derived from canonicalization techniques [27, 17, 6].
Definition 3.1.
(Orbit.) The orbit of a crystal $\mathbf{M}$ is defined as $\text{Orbit}(\mathbf{M})=\{g\cdot\mathbf{M}\mid g\in O(3)\}$.
In the context of crystal tensor prediction, all configurations of a crystal (i.e. a crystal wirh different orientations) in the space fall into the same orbit under the $O(3)$ group transformation. Specifically, for any two crystal configurations denoted as $\mathbf{M}_{1}$ and $\mathbf{M}_{2}$ within the same orbit, there exists a rotation or reflection $g\in O(3)$ such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$. Likewise, the tensor properties of the crystal also exhibit an orbit, i.e. $\text{Orbit}(\mathcal{Y})=\{g\cdot\mathcal{Y}\mid g\in O(3)\}$. All configurations of a crystal within an orbit could be transformed into one another through group transformations. Consequently, we can select a specific configuration from the orbit to serve as a canonical representative (referred to as the canonical form) that effectively characterizes the entire orbit. The formal process is outlined in the following definition.
Definition 3.2.
(Orbit Canonicalization.) A function denoted as $C_{\mathbf{M}}:\mathcal{V}\to\mathcal{V}$ is defined as an orbit canonicalization, if $\forall\mathbf{M}_{1}\in\text{Orbit}(\mathbf{M})$ and $\exists\mathbf{M}_{0}\in\text{Orbit}(\mathbf{M}),$ such that $C_{\mathbf{M}}(\mathbf{M}_{1})=\mathbf{M}_{0}$, where $\mathbf{M}_{0}$ is the canonical form of the orbit.
Orbit canonicalization enables the transformation of a crystal with arbitrary orientations in the space into one with a specific, well-defined orientation. Since the output of orbit canonicalization can always be mapped to the invariant canonical form, it follows that $C_{\mathbf{M}}(\cdot)$ itself is $O(3)$-invariant. Therefore, for any prediction function ${f_{\theta}}(\cdot)$, it holds that ${f_{\theta}}(C_{\mathbf{M}}(\cdot))$ remains $O(3)$-invariant (See the proof in Appendix A.1).
Although an $O(3)$-invariant prediction function has been established via orbit canonicalization, the complete process of equivariant prediction requires specific group transformations that map all crystal configurations within the orbit to their canonical form. To address this requirement, we introduce the concept of rigid registration. Rigid registration is a fundamental concept in fields such as computer vision and computational geometry [41]. It aims to find a transformation that matches the position and orientation of one object with the corresponding position and orientation of another object. This concept can be naturally adapted to the crystal, as follows.
Definition 3.3.
(Rigid Registration.) A function $R_{\mathbf{M}}:\mathcal{V}\times\mathcal{V}\to O(3)$ is a rigid registration if $\forall\mathbf{M}_{1},\mathbf{M}_{2}\in\text{Orbit}(\mathbf{M}),R_{\mathbf{M}}(\mathbf{M}_{1},\mathbf{M}_{2})=g$, such that $\mathbf{M}_{1}=g\cdot\mathbf{M}_{2}$.
Using Definition 3.3, we can identify a transformation that matches two configurations of a crystal. Specifically, this allows us to find specific group transformations that map all crystal configurations within the orbit to their canonical form. By leveraging Definition 3.2 and Definition 3.3 together, we derive a view for performing $O(3)$-equivariant tensor prediction through canonicalization, as described below.
Proposition 3.4.
($O(3)$-Equivariant Tensor Prediction from the Perspective of Canonicalization.) Given an arbitrary tensor prediction function $f(\mathbf{M}):\mathcal{V}\to\mathcal{W}$, we can define a new function formulated as $h(\mathbf{M})=R_{\mathbf{M}}(\mathbf{M},C_{\mathbf{M}}(\mathbf{M}))\cdot f(C_{\mathbf{M}}(\mathbf{M}))$, such that $h(\mathbf{M})$ is $O(3)$-equivariant for tensor prediction.
The proof of Proposition 3.4 is provided in Appendix A.6. Specifically, in this unified view of $O(3)$ equivariant tensor prediction, we leverage rigid registration, as defined in Definition 3.3, to find specific group transformations that map crystal configurations within the orbit to their canonical form. This canonical form is determined through orbit canonicalization, as described in Definition 3.2. Subsequently, we can directly apply these transformations to the network’s output, thereby effectively achieving the $O(3)$ equivariance. That is, under this perspective, there are no specific requirements or special treatments needed for the network design, as the $O(3)$ transformations applied to the crystal can be externally determined and simultaneously applied directly to the output. Building upon Proposition 3.4, the key challenge in achieving $O(3)$-equivariant tensor prediction lies in identifying the appropriate orbit canonicalization and rigid registration functions.
@@ 3.3 Our Proposed Framework: GoeCTP
In what follows, we demonstrate that polar decomposition can effectively serve as both orbit canonicalization and rigid registration functions for crystal tensor prediction. Building on this foundation, we will first introduce the core rotation and reflection (R&R) module of the proposed GoeCTP, which is to obtain the canonical form for a crystal with arbitrary spatial orientations using polar decomposition. Then, we will describe how the input crystal data is processed and introduce the property prediction network of GoeCTP. Finally, we will explain how proposed GoeCTP achieves equivariant tensor predictions. An overview of the GoeCTP framework is illustrated in Fig. 2.
Theorem 3.5.
@@
(Polar Decomposition [9, 12, 16].) An invertible matrix $\mathbf{L}\in\mathbb{R}^{3\times 3}$ can be uniquely decomposed into $\mathbf{L}=\mathbf{Q}\mathbf{H}$, where $\mathbf{Q}\in\mathbb{R}^{3\times 3}$ is an orthogonal matrix, $\mathbf{H}\in\mathbb{R}^{3\times 3}$ is a Hermitian positive semi-definite matrix.
In the fractional coordinate system, the $O(3)$ group transformations applied to a crystal affect only the lattice matrix $\mathbf{L}$, while fractional coordinates remain invariant, making it convenient for rotating the crystal configuration to the canonical form. Therefore, throughout this work, we adopts the fractional coordinate system to represent the crystal represented as $\mathbf{M}=(\mathbf{A},\mathbf{F},\mathbf{L})$. Then building on Theorem 3.5 , we can define both orbit canonicalization and rigid registration functions for crystal tensor prediction as follows.
Proposition 3.6.
(Orbit Canonicalization and Rigid Registration for Crystal Tensor Prediction.) By performing polar decomposition on the lattice matrix $\mathbf{L}=\mathbf{Q}\mathbf{H}$ of a crystal $\mathbf{M}$, we can define functions $f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H})$ and $f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q}$, where $f_{p1}(\cdot)$ and $f_{p2}(\cdot)$ correspond to performing polar decomposition on the lattice matrix. In this perspective, $f_{p1}(\cdot)$ serves as the orbit canonicalization function, while $f_{p2}(\cdot)$ serves as the rigid registration function. The crystal configuration $\mathbf{M}_{0}=(\mathbf{A},\mathbf{F},\mathbf{H})$ is thus identified as the canonical form.
The detailed proof of Proposition 3.6 is provided in Appendix A.7. Specifically, since the polar decomposition of a lattice matrix always exists and it is unique, any orthogonal matrix acting on the lattice matrix can be separated through polar decomposition, yielding a unique $\mathbf{H}$. Thus, $\mathbf{H}$ can serve as the canonical form. Moreover, the rationale of using polar decomposition is that it is a continuous canonicalization, which can provide improved robustness [6]. For clarity, we provide details on this continuity in Appendix A.14, and further discussions on constructing alternative canonical forms based on polar decomposition in Appendix A.15, respectively. Building upon Proposition 3.6, we utilize polar decomposition to construct the proposed GoeCTP model.
R&R Module. The primary function of the R&R Module is to rotate the crystal configuration to its canonical form and obtain the corresponding transformation. Therefore, based on Proposition 3.6, the R&R Module directly applies polar decomposition to the lattice matrix $\mathbf{L}$ to construct the rigid registration and the orbit canonicalization. Formally, the R&R Module can be expressed as:
$f_{p1}(\mathbf{M})=(\mathbf{A},\mathbf{F},\mathbf{H}),\text{ }f_{p2}(\mathbf{M},f_{p1}(\mathbf{M}))=\mathbf{Q},$ (4)
@@
where $f_{p1}(\cdot)$ is the orbit canonicalization function, while $f_{p2}(\cdot)$ is the rigid registration function. As shown in Fig. 2, the crystal configuration $(\mathbf{A},\mathbf{F},\mathbf{H})$ derived from orbit canonicalization is passed to the subsequent crystal graph construction module for further processing. The orthogonal matrix $\mathbf{Q}$ obtained during this decomposition is passed to the reverse R&R module, ensuring the equivariant transformation of the output tensor properties. Our proposed R&R module based on polar decomposition allows the input crystal data to be transformed into the canonical form that is invariant under $O(3)$ space group transformations. With this particular module, the equivariance can be captured, meaning that the subsequent components of GoeCTP are no longer required to account for equivariance.
Crystal Graph Construction. To enable networks to handle such infinite crystal structures $(\mathbf{A},\mathbf{F},\mathbf{H})$, it is typically necessary to employ graph construction methods that represent the interactions between infinite crystal structures and atoms using finite graph data. Here, we use the crystal graph construction from Yan et al. [50, 51] to describe the structure and relationships within crystals. A detailed explanation of the crystal graph construction can be found in Appendix A.2. Upon construction, the edge features are denoted as $e_{ij}$, while the node features, representing atomic properties, are denoted as $\bm{a}_{i}$.
@@
Node and Edge Feature Embedding. Building on previous work [47, 51, 50], node features $\bm{a}_{i}$ are embedded into a 92-dimensional CGCNN feature vector $\bm{f}_{i}$. Edge features $\mathbf{e}_{ij}$ are decomposed into their magnitude $||\mathbf{e}_{ij}||_{2}$ and a normalized direction vector $\hat{\mathbf{e}}_{ij}$. The magnitude is further mapped to a term similar to potential energy, $-c/||\mathbf{e}_{ij}||_{2}$, through the application of a radial basis function (RBF) kernel for encoding [21]. Subsequently, $\mathbf{e}_{ij}$ are embedded into feature vector $\bm{f}_{ij}^{e}$.
@@
Prediction Module. Since the R&R module is responsible for preserving equivariance, any predictive network can serve as the Prediction module, such as those proposed by Yan et al. [50], Taniai et al. [42], Yan et al. [49], Lee et al. [20], Wang et al. [44], among others. For better performance, we select eComformer [50], which has demonstrated excellent performance in single-value property prediction tasks, as our Prediction module. A detailed explanation of eComFormer can be found in Appendix A.2. Once processed through the stacked layers of eComFormer, the node features are aggregated to generate the crystal’s global features as follows: $\bm{G}^{\mathrm{final}}=\frac{1}{n}\sum_{1\leq i\leq n}\bm{f}_{i}^{\mathrm{final}}$.
@@
$\bm{\varepsilon}=f_{MLP}(\bm{G}^{\mathrm{final}})$ (5)
@@
$\bm{\varepsilon}^{\mathrm{final}}=\mathbf{Q}\bm{\varepsilon}\mathbf{Q}^{\top}.$ (7)
For predicting the higher-order piezoelectric and elastic tensor, the conversion process becomes more complex, see Appendix A.8 for more details.